# ESCHER: Efficient and Scalable Hypergraph Evolution Representation with Application to Triad Counting

### v1

@@ III-B Operations on ESCHER
2 for each $h_{i}$ in deleted hyperedges DelDel in parallel do
@@ 4 $node\leftarrow$ root of blockManagerblockManager
15       Mark the current nodenode as affected by adding it to AffectedAffected.
@@ 31        Add parentparent node to AffectedAffected
For a set of deleted hyperedges DelDel, each parallel thread handles one hyperedge from DelDel, marking it as affected in $O(\frac{|Del|}{p}\cdot\log|E|+\log|E|)$ time. Next, the affected nodes are processed in parallel to update their availavail values. Since each level of block manager tree requires $O(|Del|)$ work and at most $\log|E|$ levels are needed to propagate updates to the root, this step also runs in $O(\tfrac{|Del|}{p}\cdot\log|E|+\log|E|)$ time.
@@
Given a set of inserted hyperedges InsIns, one thread per hyperedge is launched to identify available memory blocks, via the block manager. Each thread $Th_{j}$ is assigned to locate the $j^{th}$ node that contains an available memory block. Because each available node can be reached through a unique path, parallel threads with distinct destinations can proceed asynchronously. Algorithm 2 guides traversal by comparing the thread ID with availavail at each node and selecting an efficient path for each thread.
@@
2for each $h_{i}$ in inserted hyperedges InsIns in parallel do
@@ 6 $count_{Del}\leftarrow$ count unique triads containing hyperedges from $Aff_{Del}$ only
8 Delete hyperedges DelDel in the hypergraph
9 Insert hyperedges InsIns in the hypergraph
@@ III-C Parallel Hypergraph Triad Count Update
Using ESCHER we develop an algorithm for updating hypergraph triads, under hyperedge insertion or deletion. Algorithm 3 begins by identifying deletion-affected regions. Step 1 processes deleted hyperedges in parallel and marks them and their 1- and 2-hop neighbors as $Aff_{Del}$. In Step 2, depending on the triad category, any state-of-the-art counting technique is applied to compute the number of triads (denoted $count_{Del}$) formed only by hyperedges in $Aff_{Del}$. Thus, $count_{Del}$ represents the triad count in the deletion-affected regions. Step 3 removes the deleted hyperedges DelDel and inserts new hyperedges InsIns. Step 4 identifies insertion-affected regions $Aff_{Ins}$ by processing InsIns in parallel. Step 5 applies the same counting technique as in Step 2 to compute triad counts in the combined affected regions $Aff_{Del}\cup Aff_{Ins}$. The updated triad count is obtained by subtracting $count_{Del}$ from the previous total count countcount and adding the newly formed triads (Step 6, Algorithm 3 Line 3).
@@ V-A Analysis of ESCHER Operations in Triad Counting
We next increased the total hypergraph size and observed the execution time. We generated random hypergraphs by varying the number of hyperedges from 2020 million to $55$ million, while setting the total number of vertices to one third of the hyperedges and a maximum cardinality of $10000$ per edge. The number of changed hyperedges is fixed at 5050K. Fig. LABEL:fig:varying_graph_size shows that the execution time increases almost linearly to the number of hyperedges, indicating good scalability.

### v2

@@ III-B Operations on ESCHER
2 for each $h_{i}$ in deleted hyperedges $Del$ in parallel do
@@ 4 $node\leftarrow$ root of blockManagerblockManager
15       Mark the current nodenode as affected by adding it to $Affected$.
@@ 31        Add parentparent node to AffectedAffected
For a set of deleted hyperedges $Del$, each parallel thread handles one hyperedge from $Del$, marking it as affected in $O(\frac{|Del|}{p}\cdot\log|E|+\log|E|)$ time. Next, the affected nodes are processed in parallel to update their availavail values. Since each level of block manager tree requires $O(|Del|)$ work and at most $\log|E|$ levels are needed to propagate updates to the root, this step also runs in $O(\tfrac{|Del|}{p}\cdot\log|E|+\log|E|)$ time.
@@
Given a set of inserted hyperedges $Ins$, one thread per hyperedge is launched to identify available memory blocks, via the block manager. Each thread $Th_{j}$ is assigned to locate the $j^{th}$ node that contains an available memory block. Because each available node can be reached through a unique path, parallel threads with distinct destinations can proceed asynchronously. Algorithm 2 guides traversal by comparing the thread ID with availavail at each node and selecting an efficient path for each thread.
@@
2for each $h_{i}$ in inserted hyperedges $Ins$ in parallel do
@@ 6 $count_{Del}\leftarrow$ count unique triads containing hyperedges from $Aff_{Del}$ only
8 Delete hyperedges $Del$ in the hypergraph
9 Insert hyperedges $Ins$ in the hypergraph
@@ III-C Parallel Hypergraph Triad Count Update
Using ESCHER we develop an algorithm for updating hypergraph triads, under hyperedge insertion or deletion. Algorithm 3 begins by identifying deletion-affected regions. Step 1 processes deleted hyperedges in parallel and marks them and their 1- and 2-hop neighbors as $Aff_{Del}$. In Step 2, depending on the triad category, any state-of-the-art counting technique is applied to compute the number of triads (denoted $count_{Del}$) formed only by hyperedges in $Aff_{Del}$. Thus, $count_{Del}$ represents the triad count in the deletion-affected regions. Step 3 removes the deleted hyperedges $Del$ and inserts new hyperedges $Ins$. Step 4 identifies insertion-affected regions $Aff_{Ins}$ by processing $Ins$ in parallel. Step 5 applies the same counting technique as in Step 2 to compute triad counts in the combined affected regions $Aff_{Del}\cup Aff_{Ins}$. The updated triad count is obtained by subtracting $count_{Del}$ from the previous total count countcount and adding the newly formed triads (Step 6, Algorithm 3 Line 3).
@@ V-A Analysis of ESCHER Operations in Triad Counting
We next increased the total hypergraph size and observed the execution time. We generated random hypergraphs by varying the number of hyperedges from $20$ million to $55$ million, while setting the total number of vertices to one third of the hyperedges and a maximum cardinality of $10000$ per edge. The number of changed hyperedges is fixed at 5050K. Fig. LABEL:fig:varying_graph_size shows that the execution time increases almost linearly to the number of hyperedges, indicating good scalability.